概均質ベクトル空間に関連する文献

以下は概均質ベクトル空間に関連する文献リスト (§1) と, それに対する簡単な解説である (§2).

このリストは, 佐藤が 1989–1990 年に作りかけたものをもとに, 短期共同研究に参加された方々からのご注意をいただいて作成した. とくに, 行者明彦氏からは共同作業と言うのがふさわしいほどの御助力をいただいたが, 誤りや不適切なコメント（おそらくあるにちがいないが）の責任は佐藤にある.

当初の作成者の視野の狭さは行者氏によって大いに正されたが, CDROM 版の Math. Rev. で検索を始めてみると, とくに表現論的な視点から概均質ベクトル空間に接近している文献についてまだまだ理解と目配りが行き届いておらず, 完備した文献表と解説を現時点で仕上げるだけの力量（と時間的余裕）がいまだ備わっていないことを痛感した. 結果としてこのリストは完全なものとは言えず, 一方, 基本文献にしぼりこんでもいないため, 中途半端で使いにくいものになってしまったことを読者にお詫びしたい. 今後, 余裕ができればぜひ改訂したいとは思っている.

このように欠陥の多いものだが, 概均質ベクトル空間に関心を持つ方々にとって多少なりとも参考になれば幸いである. [ ‘95.3.23, 佐藤文広（立教大理） ]

§1 文献表

S.Abeasis

[1] On the ring of semi-invariants of the representations of an equioriented quiver of type A_{n}, Boll. Un. Mat. Ital. A (6) 1(1982), 233–240.

D.Achab

[1] Fonction zêta d’une représentation d’algèbre de Jordan, C. R. Acad. Sc. Paris 316(1993), 977–982.

[2] Fonctions zêta des représentations des algebres de Jordan, Thèse, Université Paris VI, 1993.

A.V.Alekseevsky

[1] Component groups of centralizer for unipotent elements in semi-simple algebraic groups, Trudy Tbiliss. Mat. Inst. 62(1979), 5–27.

K.Aomoto

[1] q-analogue of b-functions and Jackson integrals, preprint.

T.Arakawa

[1] Dirichlet series related to the Eisenstein series on the Siegel upper half plane, Comment. Math. Univ. St. Pauli 27(1978), 29–42.

[2] Dirichlet series corresponding to Siegel’s modular forms, Math. Ann. 238(1978), 157–173.

[3] On automorphic forms of a quaternion unitary group of degree two, J. Fac. Sci. Univ. Tokyo, Sect.IA 28(1982), 547–566.

[4] On certain automorphic forms of Sp(1, q), Automorphic forms of several variables, Progress in Math. Vol.46, Birkhäuser(1984), 1–48.

[5] Special values of L-functions associated with the space of quadratic forms and the representation of Sp(2n, _{p}) in the space of Siegel cusp forms, Adv. Studies in pure Math. 15(1989), 465–508.

[6] Dirichlet series corresponding to Siegel’s modular forms of degree n with level N, Tôhoku Math.

J. 42(1990), 261–286. M.F.Atiyah

[1] Resolution of singularities and division of distributions, Comm. Pure Appl. Math. 23(1970),

145–150. M.F.Atiyah, H.Donnelly and I.M.Singer

[1] Eta invariants, signature defects of cusps, and values of L-functions, Ann. of Math. 118(1983), 131–177.

[2] Signature defects of cusps and values of L-functions: the nonsplit case, Addendum to: “Eta invariants, signature defects of cusps, and values of L-functions”, Ann. of Math. 119(1984), 635– 637.

B.C.Berndt and R.J.Evans

[1] The determination of Gauss sums, Bull. AMS 5(1981), 107–129. I.N.Bernstein

[1] The analytic continuation of generalized functions with respect to a parameter, Funct. Anal. Appl.

6(1972), 273–285. I.N.Bernstein and S.I.Gelfand

[1] Meromorphic property of the function P^{λ}, Funct. Anal. Appl. 3(1969),68–69. I.N.Bernstein, I.M.Gelfand and V.A.Ponomarev

[1] Coxeter functors and Gabriel’s theorem, Russ. Math. Surv. 28(1973), 17–32. B.Blind

[1] Distributions zeta à plusieurs variables associées aux algèbres de Jordan simples euclidiennes, C.

R. Acad. Sc. Paris 311(1990), 215–218. S.Bochner

[1] Group invariance of Cauchy’s formula in several variables, Ann. of Math. 45(1944), 686–707. B.D.Boe

[1] Homomorphisms between generalized Verma modules, Trans. AMS 288(1985), 791–799. N.Bopp and H.Rubenthaler

[1] Fonction zêta associée à la série principale sphérique de certains espaces symétriques. C. R. Acad. Sci. Paris, t. 310(1990), 505–508.

[2] Fonction zêta associée ‘a la série principale sphérique de certains espaces symétriques, Ann. Sci. Ec. Norm. Sup. 26(1993), 701–745.

[3] Zeta functions associated to the principal spherical series of some families of real symmetric spaces,

to appear as a preprint 1995. J.L.Brylinski

[1] Transformations canoniques, dualité projective, théorie de Lefschetz, transformations de Fourier

et sommes trigonométriques, Astérisque 140-141(1986), 3–134. J.L.Brylinski, B.Malgrange and J.L.Verdier

[1] Transformation de Fourier géométrique I, C. R. Acad. Sci. Paris 297(1983), 55–58. C.J.Bushnell and I.Reiner

[1] Functional equations for Hurwitz series and partial zeta functions of orders, J. reine angew. Math.

364(1986), 130–148. A.Chaabouni Sellami

[1] Relations entre les fonctions moyennes sur l’espace préhomogène des matrices hermitiennes, C. R. Acad. Sc. Paris 302(1986), 215–218.

[2] Relations entre les fonctions moyennes sur un espace préhomogène, Bull. Sci. Math. (2) 113(1989),

213–237. K.Chandrasekhran and Raghavan Narasimhan

[1] Functional equations with multiple gamma factors and the average order of arithmetic functions,

Ann. of Math. 76(1962), 93–136. Chen Zhijie

[1] Zeta-functions associated with prehomogeneous vector spaces and Gaussian sums, Chinese Ann. Math. Ser. A 5(1984), 755–764.

[2] A classification of irreducible prehomogeneous vector spaces over an algebraically closed field of characteristic p. I, Chinese Ann. Math. Ser. A 6(1985), 39–48.

[3] A prehomogeneous vector space of characteristic 3, in “Group theory, Beijing 1984”, Lecture Notes in Math. 1185(1986), 266–276.

[4] A classification of irreducible prehomogeneous vector spaces over an algebraically closed field of characteristic 2, I, Acta Math. Sinica 2(1986), 168–177.

[5] On the prehomogeneous vector space (GL(1) × SL(3), ⊗ (Λ_{1} + Λ_{2}), V (1) ⊗ V (7)) (p = 3), J. East-China-Norm. Univ. (1986), no. 2, 32–36.

[6] A new prehomogeneous vector space of characteristic p, Chinese Ann. Math. Ser. A 8(1987), 22–35.

[7] A classification of irreducible prehomogeneous vector spaces over an algebraically closed field of characteristic p, II, Chinese Ann. Math. Ser. A 9(1988), 10–22.

U.Christian

[1] Selberg’s zeta-, L-, and Eisenstein series, Lect. Notes in Math. No.1030, Springer(1983).

[2] Maasssche L-Reihen und eine Identitat für Gausssche Summen, Abh. Math. Sem. Univ. Hamburg 54(1984), 163–175.

[3] Eisenstein series for congruence subgroups of GL(n, ), Amer. J. Math. 107(1985), 207–240.

J.W.Cogdell

[1] Congruence zeta functions for M_{2}( ) and their associated modular forms, Math. Ann. 266(1983), 141–198.

B.Datskovski

[1] The adelic zeta function associated with the space of binary cubic forms with coefficients in a function field, Trans. Amer. Math. Soc. 299(1987), 719–745.

[2] A mean value theorem for class numbers of quadratic extensions, Contemporary Math. 143(1993), 179–242.

[3] On Dirichlet series whose coefficients are class numbers of binary quadratic forms, in “A tribute to Emil Grosswald: Number theory and related analysis”, preprint, 1993.

B.Datskovski and D.J.Wright

[1] The adelic zeta function associated to the space of binary cubic forms, Part II: Local theory, J. reine angew. Math. 367(1986), 27–75.

[2] Density of discriminants of cubic extensions, J. reine angew. Math. 386(1988), 116–138.

R.Dedekind

[1] Über die Theorie der ganzen algebraische Zahlen, Dirichlet’s Vorlesungen über Zahlentheorie, Supplement XI, 1894.

P.Deligne [1] Applications de la formule des traces aux sommes trigonométriques, SGA 4\frac{1}{2}, Lecture Notes in

Math. 569(1977), 168–232. J.Denef

[1] On the evaluation of certain p-adic intergrals, in Séminaire de théorie des nombres, Progress in Math. 59, pp.25–47, Birkhäuser 1985.

[2] The rationality of the Poincaré series associated to the p-adic points on a variety, Invent. Math. 77(1984), 1–23.

[3] p-adic semi-algebraic sets and cell decomposition, J. reine angew. Math. 369(1986), 154–166.

[4] On the degree of Igusa’s local zeta function, Amer. J. Math. 109(1987), 991–1008.

[5] Multiplicity of the poles of the Poincaré series of a p-adic subanalytic set, Séminaire de théorie des nombres de Bordeaux, Exposé 43, (1987-88)

[6] Report on Igusa’s local zeta function, Astérisque 201-202-203(1991), 359–386.

[7] Local zeta functions and Euler characteristics, Duke Math. J. 63(1991), 713–721.

[8] Degree of local zeta functions and monodromy, Compositio Math. 89(1993), 207–216. J.Denef and L.van den Dries

[1] p-Adic and real analytic sets, Annals of Math. 128(1988), 79–138. J.Denef and A.Gyoja

[1] Character sums associated to prehomogeneous vector spaces, in preparation. J.Denef and F.Loeser

[1] Weights of exponential sums, intersection cohomology, and Newton polyhedra, Invent. Math. 106(1991), 275–294.

[2] Caractéristiques d’Euler-Poincaré, fonctions zêta locales et modifications analytiques, J. Amer. Math. Soc. 5(1992), 705–720.

[3] Détermination géométrique des sommes de Selberg-Evans, Bull. Soc. Math. France 122(1994), 101–119.

[4] Regular elements and monodromy of discriminants of finite reflection groups, preprint, 1994.

[5] Polyèdres de Newton et poids de sommes exponentielles, Lecture Notes in Math. 1454(1990),

217–222. J.Denef and D.Meuser

[1] A functional equation of Igusa’s local zeta function, Amer. J. Math. 113(1991), 1135–1152. J.Denef and P.Sargos [1] Polyèdre de Newton et distibution f_{+}^{s}. I, J. d’Analyse Math. 53(1989), 201–218.

[2] Polyèdre de Newton et distibution f_{+}^{s}. II, preprint. J.Denef and S.Sperber

[1] Notes on exponential sums mod p^{n} and Newton polyhedra, preprint. J.Denef and W.Veys

[1] On the holomorphy conjecture for Igusa’s local zeta function, preprint. B.Deshommes

[1] Critères de rationalité et application à la série génératrice d’un système d’équations à coefficients

dans un corps local, J. Number Theory 22(1986), 75–114. M.Du Sautoy

[1] Finitely generated groups, p-adic analytic groups, and Poincaré series, Bull. AMS 23(1990), 121–

126. L.Ehrenpreis and T.Kawai

[1] Poisson’s summation formula and Hamburger’s theorem, Publ. Res. Inst. Math. Sci. 18(1982),

413–426. A.G.Elašvili

[1] The centralizers of nilpotent elements in the semisimple Lie algebras, Trudy Tbiliss. Mat. Inst.

46(1975), 109–132. T.Enright, R.Howe and N.Wallach

[1] A classification of unitary highest weight modules, in “Representation theory of reductive groups,

P.C.Trombi ed.”, Progress in Math. 40, 97–143, Birkhäuser, 1983. P.Epstein

[1] Zur Theorie allgemeiner Zetafunktionen I, Math. Ann. 56(1903), 615–644.

[2] Zur Theorie allgemeiner Zetafunktionen II, Math. Ann. 63(1907), 205–216. R.J.Evans

[1] Identities for products of Gauss sums over finite fields, l’Enseignement Math. 27(1981), 197–209. G.Fujisaki

[1] On the zeta functions of the simple algebra over the field of rational numbers, J. Fac. Sci. Univ. Tokyo 7(1958), 567–604.

[2] On the L functions of the simple algebra over the field of rational numbers, J. Fac. Sci. Univ.

Tokyo 9(1962), 293–311. Fulvio Ricci and E.M.Stein

[1] Homogeneous distributions on spaces of hermitian matrices, J. reine angew. Math. 368(1986),

142–164. L.Gårding

[1] Linear hyperbolic differential equations with constant coefficients, Acta Math. 85(1950), 1–62. S.S.Gelbart

[1] Fourier analysis on matrix spaces, Memoirs of Amer. Math. Soc., 1971. I.M.Gelfand

[1] Some aspects of functional analysis and algebra, Proc. Int. Congr. Math. 1954, Amsterdam 1(1957), 253–276.

[2] Collected Papers, vol. 3, part V, Springer, 1989. I.M.Gelfand and G.E.Shilov

[1] Generalized functions Vol.1, Academic Press, New York, 1964. S.G.Gindikin

[1] Cauchy’s problem for strongly homogeneous differential operators, Trudy Moskov. Mat. Obsc.

16(1967), 181–208. S.G.Gindikin and B.R.Vajnberg

[1] On a strong form of Huygens’ principle for a class of differential operators with constant coefficients,

Trudy Moskov. Mat. Obsc. 16(1967), 151–180. H.Gradl and S.Walcher

[1] On a class of inversions, Comm. Algebra 20(1992), 2371–2392. R.Godement and H.Jacquet

[1] Zeta functions of simple algebras, Lect. Notes in Math. No.260, Springer, 1972. A.Gyoja

[1] Gauss sums of prehomogeneous vector spaces, preprint.

[2] A counter example in the theory of prehomogeneous vector spaces, Proc. Japan Acad. 66(1990), 26–27.

[3] Construction of invariants, Tsukuba J. Math. 14(1990), 437–457.

[4] Representations of reductive group schemes, Tsukuba J. Math. 15(1991), 335–346.

[5] Vector valued invariants of prehomogeneous vector spaces, J. Math. Soc. Japan 43(1991), 117–131.

[6] Invariants, Nilpotent orbits, and prehomogeneous vector spaces, J. Algebra 142(1991), 210–232.

[7] Theory of prehomogeneous vector spaces without regularity condition, Publ. RIMS. 27(1991), 861–922.

[8] On the regularity of prehomogeneous vector spaces, Proc. Japan Acad. 68(1992), 341–344.

[9] Lefschetz principle in the theory of prehomogeneous vector spaces, Adv. Studies in pure Math. 21(1992), 87–99.

[10] Bernstein-Sato’s polynomial for several analytic functions, J. Math. Kyoto Univ. 33(1993), 399– 411.

[11] Local b-functions of prehomogeneous Lagrangians, J. Math. Kyoto Univ. 33(1993), 413–436.

[12] Further generalization of generalized Verma modules, Publ. RIMS Kyoto Univ. 29(1993), 349–395.

[13] Highest weight modules and b-functions of semi-invariants, Publ. RIMS Kyoto Univ. 30(1994), 353–400.

[14] A theorem of Chevalley type for prehomogeneous vector spaces, to appear in J. Math. Soc. Japan

[15] Theory of prehomogeneous vector spaces, II, preprint.

[16] Mixed Hodge theory and prehomogeneous vector spaces, preprint.

[17] 概均質ベクトル空間の理論, 数理解析研究所講究録 718(1990), 1–128.

[18] 尾関育三氏の予想について, 数理解析研究所講究録 718(1990), 129–143.

[19] 概均質ベクトル空間の理論（筑波大学での集中講義の補足） , 数理解析研究所講究録 718(1990), 144–164.

[20] 概均質ベクトル空間の最近の発展, to appear in “数学”. A.Gyoja and N.Kawanaka

[1] Gauss sums of prehomogeneous vector spaces, Proc. Japan Acad. 61(1985), 19–22.

[2] 概均質ベクトル空間のガウス和, 数理解析研究所講究録 555(1985), 32–47. K.Hashimoto

[1] The dimension of the spaces of cusp forms on Siegel upper half plane of degree two (I), J. Fac. Sci. Univ. Tokyo 30(1983), 403–488.

[2] Representation of the finite symplectic group Sp(4, _{p}) in the space of Siegel modular forms,

Comtemporary Math. 53(1986), 253–276. E.Hecke [1] Über die Zetafunktion beliebiger algebraischer Zahlkörper, Methematische Werke, 159–171.

[2] Eine neue Art von Zetafunktionen und ihre Beziehungen zur Verteilung der Primzahlen, Erste Mitteilung, Math. Z. 1(1918), 357–376.

[3] Eine neue Art von Zetafunktionen und ihre Beziehungen zur Verteilung der Primzahlen, Zweite

Mitteilung, Math. Z. 6(1920), 11–51. D.Hejhal

[1] Some Dirichlet series with coefficients related to period of automorphic eigenforms, Proc. Japan

Acad. 58(1982), 413–417. K.Hey

[1] Analytische Zahlentheorie in Systemen hyperkomplexer Zahlen, Diss. Hamburg (1929). Y.Hironaka

[1] Spherical functions of symmetric and hermitian forms I, Japanese J. Math. 14(1988), 203–223.

[2] Spherical functions of symmetric and hermitian forms II, Japanese J. Math. 15(1989), 15–51.

[3] Spherical functions of symmetric and hermitian forms III, Tôhoku Math. J. 40(1988), 651–671. Y.Hironaka and F.Sato

[1] Spherical functions and local densities of alternating forms, Amer. J. Math. 110(1988), 473–512.

[2] Local densities of alternating forms, J. Number Theory 33(1989), 32–52.

[3] Fourier-Eisenstein transform and Plancherel formula for rational binary quadratic forms, Nagoya Math. J. 128(1992), 121–151.

[4] Eisenstein series on reductive symmetric spaces and representation of Hecke algebras, J. reine

angew. Math. 445(1993), 45–108. W.Hoffmann

[1] The non-semi-simple term in the trace formula for rank one lattices, J. reine angew. Math. 379(1987), 1–21.

[2] The trace formula for Hecke operators over rank one lattice, J. Funct. Anal. 84(1989), 373–440. H.Hosokawa

[1] The Igusa local zeta function associated with the nonregular irreducible prehomogeneous vector space, Tsukuba J. Math. 15(1991), 113–119.

[2] Some results on Igusa local zeta functions associated with simple prehomogeneous vector spaces,

筑波大学学位論文 (1994.3). R.Hotta

[1] Equivariant D-modules, Proc. Wuhan CIMPA School, 1991. R.Hotta and M.Kashiwara

[1] The invariant holonomic system on a semisimple Lie algebra, Invent. Math. 75(1984), 327–358. R.Howe

[1] Perspectives on Invariant theory: Schur duality, multiplicity-free action and beyond, Israel Math.

Conference Proc. 8(1995), 1–182. R.Howe and T.Umeda

[1] The Capelli identity, the double commutant theorem and multiplicity free actions, Math. Ann.

290(1991), 565–619. T.Ibukiyama and H.Saito

[1] On zeta functions associated to symmetric matrices and an explicit conjecture on dimensions of

Siegel modular forms of general degree, Duke Math. J. Int. Math. Res. Notices 8(1992), 161–169. J.-I.Igusa

[1] Some observations on the Siegel formula, Rice University Studies, “Complex Analysis,” Proceedings 56(1970).

[2] On certain representations of semi-simple algebraic groups and the arithmetic of the corresponding invariants (1), Invent. Math. 12(1971), 62–94.

[3] A classification of spinors up to twelve, Amer. J. Math. 92(1970), 997–1028.

[4] On the arithmetic of Pfaffians, Nagoya Math. J. 47(1972), 169–198.

[5] Geometry of absolutely admissible representations, in “Number Theory, Algebraic Geometry and Commutative Algebra”, Kinokuniya, 1973, 373–452.

[6] On a certain Poisson formula, Nagoya Math. J. 53(1974), 211–233.

[7] Complex powers and asymptotic expansions I, J. reine angew. Math. 268/269(1974), 110–130.

[8] Complex powers and asymptotic expansions II, J. reine angew. Math. 278/279(1975), 307–321.

[9] Exponential sums associated with a Freudenthal quartic, J. Fac. Sci. Univ. Tokyo 24(1977), 231– 246.

[10] Some observations on higher degree characters, Amer. J. Math. 99(1977), 393–417.

[11] Lectures on forms of higher degree, Lect. Note, Tata Institute of Fundamental Researsh, Springer,1978.

[12] Some results on p-adic complex powers, Amer. J. Math. 106(1984), 1013–1032.

[13] Complex powers of irreducible algebraic curves, in “Geometry today, Roma 1984”, Progress in Math. 60, 207–230, Birkhäuser, 1985.

[14] On functional equations of complex powers, Invent. Math. 85(1986), 1–29.

[15] Some aspects of the arithmetic theory of polynomials, in “Discrete Groups in Geometry and Analysis”, Progress in Math. 67, 20–47, Birkhäuser, 1987.

[16] Some recent results on complex powers and zeta distributions, in Séminaire de théorie des nombres, Paris 1985-86, Progress in Math. 71, 67–81, Birkhäuser, 1987.

[17] Zeta distributions associated with some invariants, Amer. J. Math. 109(1987), 1–34.

[18] On a cetain class of prehomogeneous vector spaces, J. Pure Appl. Algebra 47(1987), 265–282.

[19] On the arithmetic of a singular invariant, Amer. J. Math. 110(1988), 197–233.

[20] b-Functions and p-adic integrals, Algebraic Analysis Vol.I, Academic Press, 1988, pp.231–241.

[21] Universal p-adic zeta functions and their functional equations, Amer. J. Math. 111(1989), 671–716.

[22] A problem on certain p-adic zeta functions, Israel Math. Conf. Proc. 3(1990), 67–79.

[23] Local zeta functions of certain prehomogeneous vector spaces, Amer. J. Math. 114(1992), 251–296.

[24] A stationary phase formula for p-adic integrals and its applications, Algebraic geometry and its applications, 175–194, Springer, 1994.

[25] Local zeta functions of general quadratic polynomials, Proc. Ind. Acad. (K.G.Ramanathan memorial issue) 104(1994), 177–189.

[26] 局所ゼータ関数について, 数学 46(1994), 23–38. T.Ikai

[1] ある概均質ベクトル空間の有理軌道分解, 東北大学修士論文, 1992. V.G.Kac

[1] Some remarks on nilpotent orbits, J. Algebra 64(1980), 190–213.

[2] Infinite root systems, representations of graphs and invariant theory, Invent. Math. 56(1980), 57– 92.

[3] Infinite root systems, representations of graphs and invariant theory, II J. Algebra 77(1982), 141–

162. Y.Kajima

[1] On functional equations of prehomogeneous vector spaces obtained from castling transforms. Com-

ment. Math. Univ. St. Paul. 42(1993), 49–60. S.Kasai

[1] A classification of reductive prehomogeneous vector spaces with two irreducible components I, Japan. J. Math. 14(1988), 385–418.

[2] Universal transitivity of a certain class of reductive prehomogeneous vector spaces, Tsukuba J. Math. 13(1989), 13–22.

[3] A classification of a certain class of reductive prehomogeneous vector spaces, Comm. Algebra 17(1989), 1425–1441.

[4] A classification of a certain class of reductive prehomogeneous vector spaces II, J. Algebra 129(1990), 127–135.

[5] The b-function and the holonomy diagram of a regular simple prehomogeneous vector space (GL(1)^{2}× Spin(10), half-spin rep.+vector rep.), preprint.

[6] On the microlocal structure of regular simple prehomogeneous vector spaces (GL(1)^{2} ×SL(7), Λ_{1} + Λ_{1}^{(∗)}), preprint.

[7] The b-functions of regular simple prehomogeneous vector spaces (GL(1)^{3} × SL(2m), Λ_{2} + (Λ_{1} + Λ_{1})^{(∗)}), preprint.

[8] On some double coset decompositions of the Weyl group of the simple Lie algebra of type F_{4},

preprint. S.Kasai, T.Kimura and S.Otani

[1] A classification of simple weakly spherical homogeneous spaces (I), to appear in J. of Algebra. M.Kashiwara

[1] Microlocal calculus と概均質ベクトル空間の相対不変式の Fourier 変換 (三輪哲二記), 数理解析研究所講究録 238(1975), 60–147.

[2] B-Functions and holonomic systems (Rationality of roots of b-functions), Invent. Math. 38(1976), 33–53.

[3] On the holonomic systems of linear differential equations, II, Invent. Math. 49(1978), 121–135.

[4] Vanishing cycle sheaves and holonomic systems of differential equations, Lecture Notes in Math. 1016(1983), 134–142.

[5] The universal Verma module and b-function, Advanced Studies in Pure Math. 6(1985), 67–81. M.Kashiwara and T.Kawai

[1] On the characteristic variety of a holonomic system with regular singularities, Adv. in Math. 34(1979), 163–184.

[2] On holonomic systems for ∏_{l=1}^{N}(f_{l} + \sqrt{−1}0)^{λ_{l}}, Publ. RIMS 15(1979), 551–575.

[3] Micro-local properties of ∏_{j=1}^{n} f_{j+}^{s_{j}}, Proc. Japan Acad. 51(1975), 270–272. M.Kashiwara, T.Kimura and M.Muro

[1] Microlocal calculus of simple holonomic systems and its applications, manuscript.

M.Kashiwara and P.Schapira

[1] Microlocal study of sheaves, Astérisque 128(1985), 1–235. N.M.Katz

[1] Sommes exponentielles, Astérisque 79(1980), 1–209.

[2] Gauss sums, Kloosterman sums and monodromy groups, Annals of Math. Studies, Princeton Uni-

versity Press, 1988. N.M.Katz and G.Laumon

[1] Transformation de Fourier et majoration de sommes exponentielles, Publ. IHES 62(1985), 145–202. N.Kawanaka

[1] Fourier transforms of nilpotently supported functions on a simple Lie algebra over a finite field, Invent. Math. 69(1982), 411–435.

[2] Open problem, in “Proc. of Conference on “Algebraic groups and their representations”, p.13, 1983.

[3] Generalized Gelfand-Graev representations and Ennola duality, Advanced Studies in Pure Math. 6(1985), 175–206.

[4] Generalized Gelfand-Graev representations of exceptional simple algebraic groups over a finite field I, Invent. Math. 84(1986), 575–616.

[5] Orbits and stabilizers of nilpotent elements of a graded semisimple Lie algebra, J. Fac. Sci. Univ.

Tokyo 34(1987), 573–597. T.Kimura

[1] 概均質ベクトル空間の研究, 東京大学修士論文, 1973.

[2] 概均質ベクトル空間の特異軌道と b 関数, 数理研講究録 225(1973), 262–291.

[3] 概均質ベクトル空間の理論, 数学 32(1980), 97–118.

[4] Remark on some combinatorial construction of relative invariants, Tsukuba J. Math. 5(1981), 101– 115.

[5] On the construction of some relative invariants for GL(n) (n = 6, 7, 8) by the decomposition of the Young diagrams, Lecture Notes in Math. 867(1981), 38–54.

[6] The b-functions and holonomy diagrams of irreducible regular prehomogeneous vector spaces, Nagoya Math. J. 85(1982), 1–80.

[7] A classification of prehomogeneous vector spaces of simple algebraic groups with scalar multiplications, J. Algebra, 83(1983), 72–100.

[8] 概均質ベクトル空間の理論 -分類理論を中心として-, 数理解析研究所講究録 594(1986),12–21.

[9] A classification theory of prehomogeneous vector spaces, Adv. Studies in pure Math. 14(1988), 223–256.

[10] Arithmetic calculus of Fourier transforms by Igusa local zeta functions, Trans. AMS 346(1994), 297–306.

T.Kimura and S.Kasai

[1] The orbital decomposition of some prehomogeneous vector spaces, Adv. Studies in pure Math.

6(1985), 437–480. T.Kimura, S.Kasai and H.Hosokawa

[1] Universal transitivity of simple and 2-simple prehomogeneous vector spaces, Ann. Inst. Fourier

38(1988), 11–41. T.Kimura, S.Kasai, M.Inuzuka and O.Yasukura

[1] A classification of 2-simple prehomogeneous vector spaces of type I, J. Algebra 114(1988), 369–400. T.Kimura, S.Kasai, M.Taguchi and M.Inuzuka

[1] Some P.V.-equivalence and a classification of 2-simple prehomogneous vector spaces of type II,

Trans. Amer. Math. Soc. 308(1988), 433–494. T.Kimura, S.Kasai and O.Yasukura

[1] A classification of the representations of reductive algebraic groups which admit only a finite

number of orbits, Amer. J. Math. 108(1986), 643–692. T.Kimura and T.Kogiso

[1] On adelic zeta functions of prehomogeneous vector spaces with finitely many adelic open orbits, in

“Zeta functions in geometry”, Adv. Studies in pure Math. 21(1990), 21–31. T.Kimura and M.Muro

[1] On some series of regular irreducible prehomogeneous vector spaces, Proc. Japan Acad. 55(1979),

384–389. T.Kimura and I.Ozeki

[1] On the micro local structure of a regular prehomogeneous vector space associated with Spin(10) ×

GL(3) I, Proc. Japan Acad. 58(1982), 239–242. T.Kimura, F.Sato and Xiao-wei Zhu

[1] On the poles of p-adic complex powers and the b-functions of prehomogeneous vector spaces, Amer.

J. Math. 112(1990), 423–437. T.Kimura, K.Ueda and T.Yoshigaki [1] A classification of 3-simple prehomogeneous vector spaces of nontrivial type, to appear in Japanese

J. Math. M.Kinoshita

[1] On the ζ-functions of a total matrix algebra over the field of rational numbers, J. Math. Soc. Japan

17(1965), 374–408. M.Koecher

[1] Über die Dirichlet-Reihen mit Funktionalgleichung, J. reine angew. Math. 192(1953), 1–23. T.Kogiso

[1] Simple calculation of the residues of the adelic zeta function associated with the space of binary cubic forms, J. Number Theory 51(1995), 233–248.

[2] Calculation of the adelic zeta function associated with P.V. (GL(2), M(2)) over an algebraic number field, Algebras, Groups and Geometries 11(1994), 127–144.

[3] Simple calculation of the adelic zeta function associated with the space of binary cubic forms over a function field, preprint.

[4] Calculation of the adelic zeta function associated with P.V. (GL(2), M(2)) over a function field, preprint.

[5] On adelic zeta functions of prehomogeneous vector spaces with finitely many adelic open orbits II,

preprint. T.Kogiso, T.Kimura and M.Fujinaga

[1] On functional equation of prehomogeneous zeta distributions over a local field of characteristic p,

preprint. K.Koike

[1] Relative invariants of the polynomial rings over type A_{r}, Ã_{r} quivers, Adv. in Math. 86(1991), 235–262.

[2] Relative invariants of the polynomial rings over type D_{r} quivers, preprint. T.Kondo

[1] On Gaussian sums attached to the general linear groups over finite fields, J. Math. Soc. Japan

15(1963), 245–255. A.Kurihara

[1] On the values at non-positive integers of Siegel’s zeta functions of -anisotropic quadratic forms

with signature (1, n − 1), J. Fac. Sci. Univ. Tokyo 28(1982), 567–584. E.Landau [1] Über die Anzahl der Gitterpunkte in gewissen Bereichen II, Ausgewählte Abhandlungen der Git-

terpunktelehre, Berlin, 1962. G.Laumon

[1] Transformation de Fourier, constantes d’équations fonctionelles et conjecture de Weil, Publ. IHES

65(1987), 131–210. H.Leptin

[1] Die Funktionalgleichng der Zeta-Funktion einer einfachen Algebra, Abh. Math. Sem. Univ. Ham-

burg 19(1955), 198–220. F.Loeser

[1] Arrangements d’hyperplans et sommes de Gauss, Ann. Sci. Ec. Norm. Sup. 24(1991), 379–400.

[2] Fonctions d’Igusa p-adiques et polynômes de Bernstein, Amer. J. Math. 110(1988), 1–21.

[3] Fonctions zêta locales d’Igusa à plusieurs variables, intégration dans les fibres, et discriminants, Ann. Sc. Ec. Norm. Sup. 22(1989), 435–471.

[4] Fonctions d’Igusa p-adiques, polynômes de Bernstein et polyèdres de Newton, J. reine angew.

Math. 412(1990), 75–96. F.Loeser and C.Sabbah

[1] Equations aux différences finies et déterminants d’intégrales de fonctions multiformes, Comment.

Math. Helv. 66(1991), 458–503. G.Lusztig

[1] Fourier transforms on a semisimple Lie algebra over F_{q}, Lecture Notes in Math. 1271(1987), 177– 188.

[2] Vanishing properties of cuspidal local systems, Proc. Nat. Acad. Sci. USA 91(1994), 1438–1439.

[3] Study of perverse sheaves arising from graded Lie algebras, preprint.

[4] Character sheaves, Adv. in Math. 56(1985), 193–237; II, Adv. in Math. 57(1985), 226–265; III, Adv. in Math. 57(1985), 266–315; IV, Adv. in Math. 59(1986), 1–63; V, Adv. in Math. 61(1986), 103–155.

H.Maass [1] Spherical functions and quadratic forms, J. Indian Math. Soc. 20(1956), 117–162.

[2] Über die räumliche Verteilung der Punkte in Gittern mit indefiniter Metrik, Math. Ann. 138(1959), 287–315.

[3] Zur Theorie der Harmonischen Formen, Math. Ann. 137(1959), 142–149.

[4] Siegel’s modular forms and Dirichlet series, Lect. Notes in Math. No.216, Springer, 1971.

B.Malgrange

[1] Polynômes de Bernstein-Sato et cohomologie évanescente, Astérisque 101-102(1983), 243–267. P.D.Methee

[1] Sur les distributions invariantes par le groupe des rotations de Lorentz, Comment. Math. Helv.

28(1954), 225–269. D.Meuser

[1] On the rationality of certain generating functions, Math. Ann. 256(1981), 303–310.

[2] On the poles of a local zeta function for curves, Invent. Math. 73(1983), 445–465.

[3] The meromorphic continuation of a zeta function of Weil and Igusa type, Invent. Math. 85(1986), 493–514.

[4] On a functional equation of Igusa’s local zeta function, p-Adic Analysis, Proc. Trento 1989, Lecture

Notes in Math. 1454(1990), 309–313. Y.Morita

[1] An explicit formula for the dimension of spaces of Siegel modular forms of degree two, J. Fac. Sci.

Univ. Tokyo 21(1974), 167–248. A.Mortajine

[1] Classifications des espaces préhomogènes réguliers de type paraboliques et de leurs invariants relatif, Thèse, Université de Nancy, 1988.

[2] Classifications des espaces préhomogènes réguliers de type paraboliques et de leurs invariants relatif,

Travaux en cours 40, Hermann, Paris, 1991. I.Muller

[1] Décomposition orbital des espaces préhomogènes réguliers de type parabolique commutatif et application, C.R. Acad. Sci. Paris t.303(1986), 495–498.

[2] Formes quadratiques et classification d’orbites pour une classe d’espaces préhomogènes, C.R. Acad. Sci. Paris t.312(1991), 319–322.

[3] Structure et orbites de certains espaces préhomogènes de type parabolique construits avec des

racines orthogonales, preprint, 1995. I.Muller, H.Rubenthaler and G.Schiffmann

[1] Sur la structure de certaines algèbres de Lie graduées, C.R. Acad. Sci. Paris t.297(1983), 233–235.

[2] Structure des espaces préhomogènes associés à certaines algèbres de Lie graduées, Math. Ann. 274(1986), 95–123.

D.Mumford, J.Fogarty and F.Kirwan

[1] Geometric invariant theory, third enlarged edition, Springer, 1994.

W.Müller

[1] Signature defects of cusps of Hilbert modular varieties and values of L-series at s = 1, J. Differential Geom. 20(1984), 55–119.

[2] Manifolds with cusps of rank one, Spectral theory and L^{2}-index theorem, Lect. Notes in Math. No.1244, Springer, 1987.

J.Murakami

[1] _{p} 上の概均質ベクトル空間の相対不変式の複素巾のフーリエ変換について, 数理解析研究所講究録 555(1985), 85–92.

A.Murase and T.Sugano

[1] A note on zeta functions associated with certain prehomogeneous affine spaces, Adv. Studies in pure Math. 15(1989), 415–428.

[2] Zeta functions of prehomogeneous affine spaces, Nagoya Math. J. 132(1993), 91–114.

M.Muro

[1] Some prehomogeneous vector spaces with relative invariants of degree four and the formula of the Fourier transforms, Proc. Japan Acad. 56(1980),70–74.

[2] On prehomogeneous vector spaces related to binary cubic forms I, Mem. Fac. Sci. Univ. Kochi Ser.A 1(1980), 35–57.

[3] On prehomogeneous vector spaces related to binary cubic forms II, Mem. Fac. Sci. Univ. Kochi Ser.A 2(1981), 75–99.

[4] Singular spectrums of hyperfunctions and Fourier transforms of group invariant measures on singular orbits of prehomogeneous vector spaces I, Mem. Fac. Sci. Univ. Kochi Ser.A 4(1983), 55–88.

[5] Singular spectrum of hyperfunctions and Fourier transforms of group invariant measures on singular orbits of prehomogeneous vector spaces. II, The case of indefinite quadratic forms, Mem. Fac. Sci. Kochi Univ. 5 (1984), 73–102.

[6] Correction: “Singular spectrum of hyperfunctions and Fourier transforms of group invariant measures on singular orbits of prehomogeneous vector spaces. II, The case of indefinite quadratic forms”, Mem. Fac. Sci. Kochi Univ. 6(1985), 79–80.

[7] Microlocal analysis and calculations of some relatively invariant hyperfunctions related to zeta functions associated with the vector spaces of quadratic forms, Publ. Res. Inst. Math. Sci. 22(1986), 395–463.

[8] The dimension of the space of relatively invariant hyperfunctions on regular prehomogeneous vector spaces, Proc. Japan Acad. 63(1987), 66–68.

[9] Singular invariant tempered distributions on regular prehomogeneous vector spaces, J. Funct. Anal. 76(1988), 317–345.

[10] On uniqueness of hyperfunction solutions of holonomic systems, Arkiv för Mat. 26(1988), 305–314.

[11] A note on the holonomic system of invariant hyperfunctions on a certain prehomogeneous vector space, Algebraic Analysis Vol.II, Academic Press, 1989, 493–503.

[12] On zeta functions associated with the exceptional Lie group of type E_{6}, Adv. Studies in pure Math. 15(1989), 429–463.

[13] Invariant hyperfunctions on regular prehomogeneous vector spaces of commutative parabolic type, Tôhoku Math. J. 42(1990), 163–193.

H.A. Nguyen

[1] Prehomogeneous vector space defined by a semisimple algebraic group, Bull. Amer. Math. Soc.

81(1975), 402–406. J.Oesterlé

[1] Réduction modulo p^{n} des sous-ensembles analytiques fermés de _{p}^{N}, Invent. Math. 66(1982), 325–

341. S.Ogata

[1] Special values of zeta functions associated to cusp singularities, Tôhoku Math. J. 37(1985), 367–

384. T.Ono

[1] An integral attached to a hypersurface, Amer. J. Math. 90(1968), 1224–1236. M.S.Osborne and G.Warner

[1] Multiplicities of the integrable discrete series: the case of nonuniform lattices in an -rank one semisimple group, J. Funct. Anal. 30(1978), 287–310.

[2] The Selberg trace formula I: Γ-rank one lattices, J. reine angew. Math. 324(1981), 1–113. I.Ozeki

[1] On the micro local structure of a regular prehomogeneous vector space associated with SL(5) × GL(4) I, Proc. Japan Acad. 55(1979), 37–40.

[2] On the micro local structure of a regular prehomogeneous vector space associated with GL(8), Proc. Japan Acad. 56(1980), 18–21.

[3] On the microlocal structure of the regular prehomogeneous vector space associated with SL(5) ×

GL(4), Publ. RIMS 26(1990), 539–584. Chen-Li Pan [1] A generalization of Tate’s local zeta functional equations to certain twisted Igusa local zeta func-

tions, preprint, Johns Hopkins University, 1993. J.Pas

[1] Uniform p-adic cell decomposition, preprint. M.Rais

[1] Distributions homogènes sur des espaces de matrices, Bull. Soc. Math. France 30(1972), 5–109. S.Rallis and G.Schiffmann

[1] Distributions invariantes par le groupe orthogonale, Séminaire Nancy-Strasbourg 1973-75, Lect.

Notes in Math. No.497, Springer(1975), 494–642. M.Reeder

[1] Whittaker functions, prehomogeneous vector spaces and standard representations of p-adic groups,

J. reine angew. Math. 450(1994), 83–121. B.Riemann

[1] Über die Anzahl der Primzahlen unter einer gegebenen Grösse, Gesammelte mathematische Werke,

1859, 145–153. M.Riesz

[1] L’integrale de Riemann-Liouville et le problem de Cauchy, Acta Math. 81(1948), 1–223. H.Rubenthaler

[1] Distributions bi-invariantes par SL_{n}(k), Séminaire Nancy-Strasbourg 1973-75, Lect. Notes in Math. No.497, Springer(1975), 383–493.

[2] Une série dégénérée de représentations de SL(n, ), Lect. Notes in Math. No.739, Springer(1979), 427–458.

[3] Espaces vectoriels préhomogènes, sous-groupes paraboliques et sl_{2}-triplets, Comptes rend. Acad. Sci. Paris, 290(1980), 127–129.

[4] Etude de certains sl_{2}-triplets non principaux, Preprint IRMA, Strasbourg, 1981.

[5] Classification infinitésimale des formes réelles de certains espaces préhomogènes, Comptes rend. Acad. Sci. Paris, 295(1982), 55–57.

[6] Espaces préhomogènes de type parabolique, Lect. Notes in Math. Kyoto University, 14(1982), 189–221.

[7] Espaces préhomogènes de type parabolique, Thèse, Université de Strasbourg, 1982.

[8] Construction de certaines sous-algèbres remarquables dans les algèbres de Lie semi-simples, J.of Algebra 81(1983), 268–278.

[9] Paramétrisation d’orbites dans les nappes de Dixmier admissibles, Mémoire de la Soc. Math. France (nouvelle série) (colloque du Kleebach) 15(1984), 255–275.

[10] La surjectivité de l’application moyenne pour les espaces préhomogènes, J. Funct. Anal. 60(1985), 80–94.

[11] Formes réelles des espaces préhomogènes irréductibles de type paraboliques, Ann. Inst. Fourier 36(1986), 1–38.

[12] Une classification des paires duales dans les algèbres de Lie réductives, C. R. Acad. Sci. Paris 315(1992), 645–648.

[13] Algèbres de Lie et espaces préhomogènes, “Travaux en cours”, 44, Hermann, Paris 1992.

[14] Les paires duales dans les algèbres de Lie réductives, Astérisque 219, 1994.

[15] The complete list of dual pairs in the exceptional Lie algebras, to appear as a preprint 1995. H.Rubenthaler and G.Schiffmann

[1] Opérateurs différentiels de Shimura et espaces préhomogènes, Invent. Math. 90(1987), 409–442.

[2] Triplet de Weil associé à un polynôme homogène et à un espace préhomogène, C. R. Acad. Sci. Paris 305(1987), 407–410.

[3] sl_{2}-triplet associé à un polymome homogène, J. Reine Angew. Math. 408(1990), 136–158. C.Sabbah

[1] Proximité évanescente I, Compositio Math. 62(1987), 283–328.

[2] Proximité évanescente II, Compositio Math. 64(1987), 213–241. H.Saito

[1] A generalization of Gauss sums and its applications to Siegel modular forms and L-functions associated with the vector space of quadratic forms, J. reine angew. Math. 416(1991), 91–142.

[2] On L-functions associated with the vector space of binary quadratic forms, Nagoya Math. J.

130(1993), 149–176. I.Satake

[1] Special values of zeta functions associated with self dual cones, Progress in Math. Vol.14, Birkhäuser(1981), 359–384.

[2] 数論的多様体の不変量について ( -階数 1 の場合), 数学 35(1983), 210–220.

[3] On numerical invariants of arithmetic varieties of -rank one, in Automorphic forms of several variables, Progress in Math. Vol.46, Birkhäuser(1984), 353–369.

[4] On zeta functions associated with self-dual homogeneous cones, in “Number theory and related topics, Papers presented at the Ramanujan Colloquium, Bombay 1988”, TIFR and Oxford Univ. press, 1989, 177–193.

I.Satake and J.Faraut

[1] The functional equation of zeta distributions associated with formally real Jordan algebras, Tôhoku Math. J. 36(1984), 469–482.

I.Satake and S.Ogata

[1] Zeta functions associated to cones and their special values, Adv. Studies in pure Math. 15(1989), 1–27.

F.Sato

[1] Zeta functions in several variables associated with prehomogeneous vector spaces I: Functional equations, Tôhoku Math. J. 34(1982), 437–483.

[2] Zeta functions in several variables associated with prehomogeneous vector spaces II: A convergence criterion, Tôhoku Math. J. 35(1983), 77–99.

[3] Zeta functions in several variables associated with prehomogeneous vector spaces III: Eisenstein series for indefinite quadratic forms, Ann. of Math. 116(1982), 177–212.

[4] On zeta functions of ternary zero forms, J. Fac. Sci. Univ. Tokyo 28(1982), 585–604.

[5] Eisenstein series on semisimple symmetric spaces of Chevalley groups, Adv. Studies in pure Math. 7(1985), 295–332.

[6] p 進体上の概均質ベクトル空間, 筑波大集中講義記録 (1987.11)

[7] Eisenstein series on the Siegel half space of signature (1,1), Comment. Math. Univ. St. Pauli, 37(1988), 99–125.

[8] The Hamburger theorem for Epstein zeta functions, Algebraic Analysis Vol.II, Academic Press, 1989, 789–807.

[9] On functional equations of zeta distributions, Adv. Studies in pure Math. 15(1989), 465–508.

[10] Zeta functions with polynomial coefficients associated with prehomogeneous vector spaces. Preprint, 1989.

[11] The Maass zeta functions attached to positive definite quadratic forms. Adv. Studies in pure Math. 21(1992), 409–443.

[12] On the stability of branching coefficients of rational representations of reductive groups, Comment. Math. Univ. St. Pauli. 42(1993), 189–207.

[13] Introduction to zeta functions of prehomogeneous vector spaces, “Topics in Number Theory and Algebra” ed. by D.S.Kim, Proc. of Workshops in pure Math. Vol. 13, Part I, pp.1–21, The Korean Academic Council, 1993.

[14] Zeta functions of prehomogeneous vector spaces with coefficients related to periods of automorphic forms, Proc. Ind. Acad. (K.G.Ramanathan memorial issue) 104(1994), 99–135.

[15] Eisenstein series on weakly spherical homogeneous spaces and zeta functions of prehomogeneous vector spaces, Comment. Math. Univ. St. Pauli 44(1995), 129–150.

[16] Eisenstein series on weakly spherical homogeneous spaces of GL(n), preprint, 1995.

[17] Eisenstein series on Spin_{10}\GL_{16}, in preparation. F.Sato and H.Ochiai

[1] Castling transforms of prehomogeneous vector spaces and functional equations, Comment. Math.

Univ. St. Pauli 40(1991), 61–82. M.Sato

[1] 佐藤幹夫氏より一言御挨拶があります, 数学の歩み 15(1970), 1–8.

[2] 概均質ベクトル空間の理論 (新谷卓郎記), 数学の歩み 15(1970), 85–157.

[3] Theory of prehomogeneous vector spaces (Algebraic part) - The English translation of Sato’s lecture

from Shintani’s Note (translated by M.Muro), Nagoya Math. J. 120(1990), 1–34. M.Sato, M.Kashiwara, T.Kimura and T.Oshima

[1] Micro-local analysis of prehomogeneous vector spaces, Invent. Math. 62(1980), 117–179. M.Sato and T.Kimura

[1] A classification of irreducible prehomogeneous vector spaces and their invariants, Nagoya Math. J.

65(1977), 1–155. M.Sato and T.Shintani

[1] On zeta functions associated with prehomogenous vector spaces, Ann. of Math. 100(1974), 131–

170. A.Selberg

[1] Discontinuous groups and harmonic analysis, Proc. Intern. Congr. Math. Stockholm, 1962, 177–

189. F.J.Servedio

[1] Prehomogeneous vector spaces and varieties, Trans. Amer. Math. Soc. 176(1973), 421–444.

[2] Dense orbits in Z(P) the singular hypersurface of a prehomogenous vector space, J. pure appl. Algebra 10(1977), 169–175.

[3] Affine open orbits, reductive isotropy groups, and dominant gradient morphism; a theorem of Mikio

Sato, Pacific J. Math. 72(1977), 537–545. G.Shimura [1] On differential operators attached to certain representations of classical groups, Invent. Math.

77(1984), 463–488. T.Shintani

[1] On Dirichlet series whose coefficients are class numbers of integral binary cubic forms, J. Math. Soc. Japan 24(1972), 132–188.

[2] On zeta functions associated with the vector space of quadratic forms, J. Fac. Sci. Univ. Tokyo 22(1975), 25–65.

[3] On evaluation of zeta functions of totally real algebraic number fields at non-positive integers, J.

Fac. Sci. Univ. Tokyo 23(1976), 393–417. T.Shoji

[1] Geometry of orbits and Springer correspondence, Astérisque 168(1988), 61–140.

[2] 有限 Chevalley 群の既約指標, 数学 47(1995), 241–255. C.L.Siegel

[1] Über die analytische Theorie der quadratischen Formen II, Ann. of Math. 37(1936), 230–263.

[2] Über die Zetafunktionen indefiniter quadratischer Formen, Math. Z. 43(1938), 682–708.

[3] Über die Zetafunktionen indefiniter quadratischer Formen II, Math. Z. 44(1939), 398–426.

[4] Lectures on advanced analytic number theory, Lecture note, Tata Institute of Fundamental Re-

search, Bombay, 1961. T.A.Springer and R.Steinberg

[1] Conjugacy classes, in “Seminar on algebraic groups and related finite groups”, Lecture Notes in

Math. 131(1970). H.M.Stark

[1] L-functions and character sums for quadratic forms (I), Acta Arith. 16(1968), 35–50. E.M.Stein

[1] Analysis in matrix spaces and some new representations on SL(n, ), Ann. of Math. 86(1967),

461–490. L.Strauss

[1] Poles of a two-variable p-adic complex power Trans. AMS 278(1983), 481–493. S.Suga

[1] Highest weight modules associated with classical irreducible regular prehomogeneous vector spaces of commutative parabolic type, Osaka J. Math. 28(1991), 323–346.

T.Suzuki [1] 概均質ベクトル空間の相対不変式の Fourier 変換について, 名古屋大学修士論文, 1975.

[2] On zeta functions associated with quadratic forms of variable coefficients, Nagoya Math. J. 73(1979), 117–147.

[3] Distribution with automorphy and Dirichlet series, Nagoya Math. J. 73(1979), 157–169. T.Tamagawa

[1] On the ζ-functions of a division algebra, Ann of Math. 77(1963), 387–405. J.Tate

[1] Fourier analysis in number fields and Hecke’s zeta functions, Algebraic number theory (ed. by

J.W.S.Cassels and A.Fröhlich), Academic Press, 1967, 305–347. A.Tengstrand

[1] Distributions invariant under an orthogonal group of arbitrary signature, Math. Scand. 8(1960),

201–218. Y.Teranishi

[1] Relative invariants and b-functions of prehomogeneous vector spaces (G×GL(d_{1},…, d_{r}), ρ̃_{1}, M(n, )), Nagoya Math. J. 98(1985), 139–156.

[2] The functional equation of zeta distributions associated with prehomogeneous vector spaces (G̃, ρ̃, M(n, )),

Nagoya Math. J. 99(1985), 131–146. Liang Chi Tsao

[1] Exponential sums over finite simple Jordan algebras and finite simple associative algebras, Duke

Math. J. 42(1975), 333–345. T.Umeda

[1] 不変式論・入門・以前, 論集「現代の母関数」 , 1991, pp.71–188.

[2] 100 年目の Capelli Identity, 数学 46(1994), 206–227. W.Veys

[1] On the poles of Igusa’s local zeta function for curves, J. London Math. Soc. 41(1990), 27–32.

[2] Relations between numerical data of an embedded resolution, Amer. J. Math. 113(1991), 573–592.

[3] Congruences for numerical data of an embedded resolution, Compositio Math. 80(1991), 151–169.

[4] Numerical data of resolutions of singularities and Igusa’s local zeta function, Dissertation, Katholieke Univ. Leuven, 1991.

[5] Holomorphy of local zeta functions for curves, Math. Ann. 295(1993), 635–641.

[6] Poles of Igusa’s locla zeta function and monodromy, Bull. Soc. Math. France 121(1993), 545–598.

[7] On Euler characteristics associated to exceptional divisors, preprint.

[8] Determination of the poles of the topological zeta function for curves, preprint. G.Warner

[1] Selberg’s trace formula for non-uniform lattices: The -rank one case, Advances in Math. Supple-

mentary Studies 6(1979), 1–142. A.Weil

[1] Adeles and algebraic groups, Institute for Advanced Study, Princeton, N.J., 1961.

[2] Sur la formule de Siegel dans la théorie des groupes classiques, Acta. Math. 113(1965), 1–87.

[3] Fonction zêta et distributions, Collected papers Vol.III, 158–163.

[4] On Eisenstein’s copy of the Disquisitiones, in “Algebraic Number Theory - in honor of K.Iwasawa”, Adv. Studies in pure Math. 17(1989), 463–469.

[5] Prehistory of the zeta function, in “Number theory, Trace formulas and discrete groups (Symposium

in honor of Atle Selberg), Academic Press, 1989, 1–9. R.Weissauer

[1] Siegel modular forms and Dirichlet series, preprint, 1986. D.J.Wright

[1] The adelic zeta function associated with the space of binary cubic forms I: Global theory, Math. Ann. 270(1985), 503–534.

[2] Cubic character sums of cubic polynomials, Proc. Amer. Math. Soc. 100(1987), 409–413.

[3] Distribution of discriminants of abelian extensions, Mathematica Göttingensis Heft 38, 1987.

[4] Twists of the Iwasawa-Tate zeta function, Mathematica Göttingensis Heft 39, 1987. D.J.Wright and A.Yukie

[1] Prehomogeneous vector spaces and field extensions, Invent. Math. 110(1992), 283–314. H.Yamada

[1] Relative invariants of prehomogeneous vector spaces and a realization of certain unitary representations I, Hiroshima Math. J. 11(1981), 97–109.

[2] The b-functions and unitary representations, preprint. T.Yano [1] On the complete localization of prehomogeneous vector spaces, Saitama Math. J. 5(1987), 27–33.

[2] Locally prehomogeneous spaces and their transverse localizations, Algebraic Analysis Vol.II, Academic Press, 1989, 927–942.

[3] Poles of Igusa local zeta functions for certain class of polynomials, 代数幾何学シンポジウム（埼玉

大学）報告集, 1993, pp.185–197. T.Yano and J.Sekiguchi

[1] The micro local structure of weighted homogeneous polynomials associated with Coxeter systems

I, II, Tokyo J. Math. 2(1979), 193–220; 4(1981), 1–34. Kefeng Ying

[1] On the convergence of the adelic zeta functions associated to irreducible regular prehomogeneous vector spaces, to appear in Amer. J. Math.

[2] Numbers of solutions of some invariants in finite fields, preprint, MSRI, 1994.

[3] On the generalized global zeta functions associated to irreducible regular prehomogeneous vector

spaces, preprint, MSRI, 1994. A.Yoshimoto

[1] On a generalization of Hamburger’s theorem, Nagoya Math. J. 98(1985), 67–76. A.Yukie

[1] On Shintani zeta function associated to the space of binary quadratic forms, Math. Ann. 292(1992), 355–374.

[2] Shintani zeta functions, Lect. Notes of London Math. Soc. No.183, 1993.

[3] Prehomogeneous vector spaces, Eisenstein series and invariant theory, Mathematica Göttingensis.

[4] On Shintani zeta function for GL(2), Preprint.

[5] On the convergence of zeta functions for certain prehomogeneous vector spaces, Preprint.

[6] On parabolic D_{4}-type prehomogeneous vector space, in preparation. X.W.Zhu

[1] The classification of spinors under GSpin_{14} over finite fields, Trans. Amer. Math. Soc. 333(1992),

95–114. M.Zorn

[1] Note zur analytischen hyperkomplexen Zahlentheorie, Abh. Math. Sem. Univ. Hamburg 9(1933), 197–201.

§2 解説

2.1 概均質ベクトル空間前史

概均質ベクトル空間の理論は, 1960 年前後に佐藤幹夫氏によって始められた. 当時の佐藤幹夫氏自身の問題意識は M.Sato [1] に述べられている. その後の発展の過程については T.Kimura [8] で振り返られている. （発展の過程に興味がある場合は, 数理研講究録に現れた論説を順に眺めるとよい. とくに, No. 225, 238, 260, 416, 497, 555, 629, 718 が, b-関数および概均質ベクトル空間を中心にしてまとめられている. ）

さて, M.Sato [1] によると概均質ベクトル空間の概念を見いだした動機として, 基本解が explicit に与えられるような定数係数偏微分作用素の族を求めることがあった. この問題と多項式の複素巾の Fourier 変換, 群の作用の下での不変性との関連については, M.Riesz [1], L.Gårding [1], S.Bochner [1], I.M.Gelfand [1], I.M.Gelfand and G.E.Shilov [1], S.G.Gindikin [1], S.G.Gindikin and B.R.Vajnberg [1] を挙げておく. 一方 Theta 変換公式に基いてその関数等式が証明される多くの古典的ゼータ関数が存在した（例えば Riemann ゼータ関数： B.Riemann [1] （ A.Weil [4], [5] も参照せよ） , Epstein ゼータ関数： P.Epstein [1], [2], Dedekind ゼータ関数： E.Hecke [1], 単純環のゼータ関数： K.Hey [1], M.Zorn [1], H.Leptin [1], G.Fujisaki [1], [2], 不定値二次形式のゼータ関数： C.L.Siegel [2], [3], det^{t}xx (x は m × n 行列) のゼータ関数： M.Koecher [1] ） . これらを統一的に見る視点は, 一つには保型形式の理論によって与えられることになるが, 概均質ベクトル空間のゼータ関数の理論もまた別の視点を提供することになる.

2.2 代数的理論

概均質ベクトル空間の代数的理論（相対不変式, 正則性, 等）は M.Sato [2], [3], M.Sato and T.Kimura [1], A.Gyoja [7] にまとめられている. 定義体を考慮に入れた議論は F.Sato [1] で多少なされている.

2.3 分類

既約概均質ベクトル空間の分類は M.Sato and T.Kimura [1] で行われた. 次の課題は既約でない reductive 概均質ベクトル空間の分類であり, 木村達雄氏を中心とする Tsukuba school で精力的に研究が進められている. 現在のところ

(1) 軌道の個数が有限個の場合： T.Kimura, S.Kasai and O.Yasukura [1];

(2) 作用する群の単純成分の個数が < 3 の場合： T.Kimura [7], T.Kimura, S.Kasai, M.Inuzuka and O.Yasukura [1], T.Kimura, S.Kasai, M.Taguchi and M.Inuzuka [1];

(3) 作用する群の単純成分の個数が 3 の場合： T.Kimura, K.Ueda and T.Yoshigaki [1];

(4) 既約直和因子の個数が 2 の場合： S.Kasai [1];

(5) いわゆる trivial 概均質ベクトル空間を直和因子に含まない場合： S.Kasai [3];

(6) 既約直和因子がすべて正則概均質ベクトル空間の場合： S.Kasai [4];

などの場合に分類がなされている. これらの研究を通じて, 分類にとっての躓きの石が trivial 概均質ベクトル空間にあることが明らかになったように思われる. trivial 概均質ベクトル空間については, quiver との関係が注目されるべきであろう（たとえば, V.G.Kac [2], [3], K.Koike [1], [2] ） . 分類論の現状の総合報告として T.Kimura [8], [9] があったが, いささか古くなってしまった.

以上は標数 0 の代数閉体上での分類であるが, 実数体上の分類としては, 既約正則かつ被約の場合に T.Kimura による未発表の結果がある (T.Suzuki [1] 参照). H.Rubenthaler [5], [11] は既約正則放物型の概均質ベクトル空間（既約正則被約ということとほとんど一致する）に対して実数体上の分類を行っている.

概均質ベクトル空間に付随するゼータ関数がオイラー積を持つことと密接に関係する条件として「普遍推移性」 (universal transitivity) があるが, J.Igusa [14], [18] は既約正則な概均質ベクトル空間の split form に対して「普遍推移性」をもつものを分類した. S.Kasai [2], T.Kimura, S.Kasai and H.Hosokawa [1] は, 可約な概均質ベクトル空間に対して同じ問題を調べている.

標数 p の体上での分類は Z.Chen [2]–[7] が研究している.

2.4 b-関数

概均質ベクトル空間の b-関数の一般論は, M.Sato [2], [3], M.Sato, M.Kashiwara, T.Kimura and T.Oshima

[1] が基本的である. 超局所 b-関数（１変数）については, その計算法が M.Sato, M.Kashiwara, T.Kimura and T.Oshima [1] に与えられている. （この論文では, アルゴリズムの一部が欠落していたが A.Gyoja [13, 5.21] で補充された. ）超局所 b-関数（多変数）については, M.Kashiwara, T.Kimura and M.Muro [1], A.Gyoja [11], [13] を見よ.

M.Sato, M.Kashiwara, T.Kimura and T.Oshima [1] の方法を用いて, 既約概均質ベクトル空間の b-関数は explicit に決定されている (T.Kimura [6], T.Kimura and M.Muro [1], T.Kimura and I.Ozeki [1], I.Ozeki [1], [2], A.Gyoja [7, 3.24], T.Yano and I.Ozeki (unpublished)). 既約でない空間の b-関数は Y.Teranishi [1], M.Muro [9], F.Sato [11], S.Kasai [5], [6], [7], などで調べられている. （相対不変式の複素巾の Fourier 変換の公式（上の関数等式）が explicit に与えられれば, それからも b-関数は分かるので, 「ゼータ関数の関数等式」の (i) 項も参照せよ. ）

b-関数の零点は相対不変式の複素巾の極の位置を記述しているが, Igusa は, 相対不変式の p-進複素巾についてもその極の位置は b-関数の零点によって統制されることを多数の具体例の計算を通じて確認した (J.Igusa [20] 参照). 既約正則概均質ベクトル空間においてこのことが一般的に成立することは, T.Kimura, F.Sato and Xiao-wei Zhu [1] で（分類に依拠する点が不満だが）示された（ T.Yano [3] に訂正点が指摘されている） .

b-関数は, 群の多項式環上の表現に関連させる形で拡張することができる. これについては G.Shimura [1], M.Muro (unpublished work), H.Rubenthaler and G.Schiffmann [1], F.Sato [10], [11] がある.

相対不変式と限らぬ一般の多項式の複素巾の解析接続については I.N.Bernstein and S.I.Gelfand [1], M.F.Atiyah [1] で特異点解消を用いて示された. p-進複素巾の解析接続は J.Igusa [7], [8] で示された. Igusa 氏によるこの辺りの仕事は J.Igusa [11] にまとめて解説されている. I.N.Bernstein and S.I.Gelfand [1] には, 特異点解消を用いる方法が p-進体にも適用されることが示唆されていることも付記しておく. 一般の多項式に対する b-関数の存在とそれに基づく複素巾の解析接続の新しい証明は I.N.Bernstein [1] で与えられた.

多項式の複素巾を D-加群の理論の側から研究したものとして M.Kashiwara [2], [3], [4], M.Kashiwara and T.Kawai [1], A.Gyoja [7], [15] がある. これをエタール層の理論の側から研究したものとして P.Deligne [1], N.M.Katz [1], N.M.Katz and G.Laumon [1], G.Laumon [1], F.Loeser [1], J.Denef and F.Loeser [3] がある.

b-関数の零点が負の有理数になることは, M.Kashiwara [2] で証明された. この結果は, 多変数の b-関数に一般化される. C.Sabbah [2], A.Gyoja [10] を見よ.

b-関数と vanishing cycle との関係については B.Malgrange [1], M.Kashiwara [4] を見よ. この関係は概均質ベクトル空間の研究に層の理論が関わってくるときには, 常に基本的である. J.Denef and A.Gyoja [1], A.Gyoja [16] を見よ.

geometric Fourier 変換については, J.L.Brylinski [1], J.L.Brylinski, B.Malgrange and J.L.Verdier [1], R.Hotta and M.Kashiwara [1], M.Kashiwara and P.Schapira [1], N.M.Katz and G.Laumon [1] を見よ.

2.5 ゼータ関数の関数等式

概均質ベクトル空間の相対不変式の複素巾の（超関数としての） Fourier 変換が, 双対的な概均質ベクトル空間の相対不変式の複素巾で表されるという事実が概均質ベクトル空間のゼータ関数の関数等式である. ゼータ関数, 及びその関数等式は, アルキメデス的局所体（ , ） , 非アルキメデス的局所体, 有限体, 代数的数体上で考えることができる.

(i) アルキメデス的局所体（ , ）この場合は, まず作用する群が reductive , 特異点集合が超曲面で -既約成分は絶対既約の条件の下で M.Sato [2] で関数等式が示された. T.Shintani [1] には同じ証明が, 特異点集合が絶対既約というより強い条件の下で詳述されている. また同じ絶対既約の条件下で別証明が M.Sato and T.Shintani [1] にある. 一般に reductive と限らぬ概均質ベクトル空間で特異点集合が超曲面となるものに対しては F.Sato [1] で正則部分空間に関する部分 Fourier 変換に拡張された形で述べられている. さらに A.Gyoja [7] では, 作用する群が reductive との仮定の下で, 正則とは限らぬ概均質ベクトル空間に対しても関数等式が示されている.

関数等式の具体的な形の決定には M.Kashiwara [1] で解説されている Micro-local calculus による方法が強力なアルゴリズムを与えている (T.Kimura [3] も参照せよ). この方法に基づく計算の実行は T.Suzuki [1], M.Muro [1], [2], [3], [7], [11], [12] 等でなされた. Y.Teranishi [2], I.Satake and J.Faraut [1], I.Muller

[1] でも一連の空間に対して関数等式の具体的な決定を行っている. M.Sato [2], T.Shintani [1], [2], T.Suzuki [2], F.Sato [1], [3], [9], [11], [16], [17], B.Datskovski and D.J.Wright [1] にも関数等式の具体例が含まれている. 二次形式の場合は I.M.Gelfand and G.E.Shilov [1], S.Rallis and G.Schiffmann [1], 全行列環の行列式の場合は E.M.Stein [1], S.S.Gelbart [1], 単純環のノルム形式の場合は表現付きの一般の形で R.Godement and H.Jacquet [1] で扱われている.

関数等式の表現付（球関数付）ゼータ関数への拡張は, H.Rubenthaler and G.Schiffmann [1], F.Sato [10] (有限次元表現の場合), N.Bopp and H.Rubenthaler [1], [2], [3], F.Sato [14] (無限次元表現の場合) で扱われている.

(ii) 非アルキメデス的局所体標数 0 の非アルキメデス的局所体, すなわち p-進体の場合, 相対不変式の p-進複素巾の Fourier 変換に関する関数等式は, 作用する群が self-adjoint, 特異点集合が絶対既約超曲面のときに軌道の個数の有限性を仮定し J.Igusa [12] で証明された. この結果は必ずしも reductive でない正則概均質ベクトル空間の多変数ゼータ関数にも軌道の有限性の仮定を適当な形で課するならば拡張することができる（ F.Sato [7], [8]). さらに, 軌道の個数の有限性という条件をはずすことは, reductive 概均質ベクトル空間に対して A.Gyoja (unpublished work) によってなされた.

p-進体上の局所ゼータ関数が p^{−s} の有理関数になるという重要な事実は, J.Igusa [8] で証明され, さらに J.Denef [1], [2], B.Deshommes [1] などで拡張されている. p-進局所ゼータ関数については, J.Igusa [26], J.Denef [6] という理論発展の中心人物による良い survey があるので参照されたい.

p-進体上の関数等式の具体例は J.Tate [1] を別とすれば, まず二次形式の場合が S.Rallis and G.Schiffmann [1] で与えられた. 単純環のノルム形式についてはやはり R.Godement and H.Jacquet [1] を見よ. J.Igusa [14], [17], I.Muller [1], F.Sato [6], [9] も p-進体上の関数等式の具体例を扱っている. このうち J.Igusa [14], I.Muller [1], F.Sato [9] では非アルキメデス的局所体とアルキメデス的局所体を並行させて議論している. Y.Hironaka and F.Sato [2] では, 交代行列の local density の計算に p-進複素巾の関数等式を応用した. Y.Hironaka [1], [2], [3], Y.Hironaka and F.Sato [1] で調べられている球関数はある概均質ベクトル空間に付随する p-進局所ゼータ関数ともみなすことができる.

(iii) 有限体有限体の場合, 相対不変式の複素巾に対応しているのは相対不変式と有限体の乗法群の指標の合成であり, 関数等式はその Fourier 変換がある定数倍を除いて再び相対不変式と指標の合成に等しいという形をとると期待される. この際現れる定数は Gauss 和の一般化とみなせる. このことは N.Kawanaka によって予想され, A.Gyoja and N.Kawanaka [1], A.Gyoja [1] において既約正則概均質ベクトル空間に対して証明された.

最近, 有限体上の概均質ベクトル空間の関数等式の具体的な形は, 標数が十分大きいとの仮定のもとで J.Denef and A.Gyoja [1] により決定された. 個別例については, それ以前に, 少なくとも次の場合には決定されていた. 番号は M.Sato and T.Kimura [1] の既約概均質ベクトル空間の分類の表に従う. (I) 正則の場合：二次形式 H.M.Stark [1]. (1), (2), (3), (13), (15) Z.Chen [1]. （ J.Murakami [1] も見よ. ） (4) R.Evans [1] の議論を参考に計算可能. (27) L.Tsao [1]. (II) 非正則の場合： A.Gyoja [17]. また球関数つきのゼータ関数の有限体上の類似に関しては (1) T.Kondo [1], (3) A.Gyoja [5] を見よ.

有限体上の関数等式の応用には,

(1) 有限 reductive 代数群のある種の既約表現の指標の unipotent 元での値の記述（ N.Kawanaka [4] ） , これが Kawanaka-Gyoja 両氏の研究のそもそもの動機であった, （「表現論との関係」の項参照）

(2) 数論的応用として概均質ベクトル空間に付随する L-関数の関数等式 (H.M.Stark [1], F.Sato [9], B.Datskovski and D.J.Wright [1], T.Arakawa [5], H.Saito [1], [2]), Siegel 保型形式の指標による twist (R.Weissauer [1]) などがある.

(iv) 代数的数体 M.Sato [2] では, 関数等式を満たす超関数値ゼータ関数が構成されている. Dirichlet 級数としてのゼータ関数は T.Shintani [1] における具体的な場合（ 2 元 3 次形式の空間）の考察を経て M.Sato and T.Shintani [1] で一変数ゼータ関数の一般論が築かれた. T.Shintani [2] では対称行列の空間のゼータ関数が詳しく調べられている. T.Shintani [1] の結果は, D.Wright [1], B.Datskovski and D.J.Wright [1], [2], B.Datskovski

[1] で adelic language を用いて一般の大域体に拡張された. 多変数ゼータ関数の理論は, T.Shintani [2], T.Suzuki [1], F.Sato [3] により二次形式に関連した具体例の研究がまず行われ, 一般論は F.Sato [1], [2] で与えられた.

ゼータ関数の研究の前提としては, その収束の証明がある. 個々の例に対する case by case の証明ではなく, ある程度の一般性をもった収束判定条件としては次のようなものがある. まず, A.Weil [2], J.Igusa

[2] で与えられたテータ級数の収束判定条件がある. この条件は M.Sato and T.Shintani [1] において基本的な仮定として用いられた. Weil-Igusa の条件はかなりきついもので, この条件を仮定するとゼータ関数の収束・関数等式・解析接続の証明のみならず, 極における留数まで（原理的には）計算ができる. より広いクラスに適用できる条件は F.Sato [2] で与えられた. その判定条件は多変数ゼータ関数にも適用でき, 収束域も best possible と思われるものを与える. 同じ判定条件は, 既約正則の場合に K.Ying [1] によって再発見され, 代数体上での定式化で F.Sato [2] とは若干異なる証明が与えられた. K.Ying [3] はその拡張である. 一方, Sato-Ying の条件の適用できない場合をカバーする条件として, A.Yukie [2] で与えられた dim G = dim V がある. Yukie の方法は, Weil [2] の方法の発展であり, dim G = dim V を越えた適用範囲を持つ（ A.Yukie [2], [5] 参照） . これらの研究の結果, 既約正則の場合に限ると 3 系列を除いて収束の問題は解決し, 残る 3 系列のうちの 2 系列についても被約な split form については収束が確認されている.

ゼータ関数の極における留数, より一般に特異部の決定は, 整数論への応用のためにもきわめて重要である（「数論的量の漸近評価」の項参照）のだが, Weil-Igusa の条件が満たされないときにはたいへん難しい問題であり, きちんと計算しきれている場合は多くない. T.Shintani [1], [2], T.Arakawa [2], [6], F.Sato [4] などに計算例があったが, 最近 A.Yukie [2] は T.Shintani [1] における Eisenstein 級数を用いる方法を大いに拡張して成果を挙げている.

(v) ゼータ関数の一般化ゼータ関数の理論の拡張としては次のようなものがある.

(1) 概均質ベクトル空間の L-関数 ⋯ 「有限体」の項参照.

(2) 調和多項式係数のゼータ関数 ⋯ P.Epstein [1], [2], C.L.Siegel [4], H.Maass [1], [4] 等で扱われている球関数付きゼータ関数の理論を概均質ベクトル空間の立場から整理することにより, G_{}がコンパクト因子を含むときにその一般化が得られる (F.Sato [8], [10], [11]). より一般には, G_{}の保型形式の周期を係数とするゼータ関数を考えるべきである. すなわち, H.Maass [2], D.Hejhal [1] で研究されたタイプのゼータ関数を概均質ベクトル空間の立場から調べることで関数等式の根拠を明白にし, かつ拡張することができる (F.Sato [14]). 保型形式付ゼータ関数を考えることで, 概均質ベクトル空間のゼータ関数に対する Hamburger 型の定理を示せる (L.Ehrenpreis and T.Kawai [1], A.Yoshimoto [1], F.Sato [8]).

(3) 概均質アフィン空間のゼータ関数 ⋯ A.Murase and T.Sugano [1], [2] は, 代数群がベクトル空間にアフィン変換として作用する場合にゼータ関数の理論を拡張した. こうして得られるものの中には正則でない概均質ベクトル空間のゼータ関数と見なせるものが含まれている. Murase-Sugano のこの研究は, Jacobi 形式の空間の次元公式に現われるゼータ関数を理解することを出発点としてなされた（「ゼータ関数の特殊値と保型形式の次元公式」の項参照） .

(4) その他 ⋯ reductive 等質空間に放物型部分群が概均質に作用するとき（弱球等質空間という） , 概均質ベクトル空間の場合と同様にゼータ関数（ Eisenstein 級数）を定義することができる. そのゼータ関数は, 若干の付加的条件の下で Weyl 群の作用に関して関数等式を持つなど良い性質を持つことが期待される. 対称空間の場合は F.Sato [5], [7], Y.Hironaka and F.Sato [4] で, 群は GL(n) とするが対称空間とは限らぬ場合は F.Sato [15], [16], [17] で調べられている. 群が GL(n) のとき, 弱球等質空間と概均質ベクトル空間とは密接に関連しており, 概均質ベクトル空間への新しい視角を与えると思われる.

2.6 裏返し変換

M.Sato and T.Kimura [1] の分類理論において, 裏返し変換が基本的に重要であったように, より一般の概均質ベクトル空間の全体像を理解するためには, 「裏返し変換」というものの, より深い理解が不可欠である. 裏返し変換の一般化の試みとして T.Teranishi [1], [2], I.N.Bernstein, I.M.Gelfand and V.A.Ponomarev [1], V.G.Kac [2] がある. また A.Gyoja [14], Theorem A で与えられた変換も, そのような一般化とみなすことができるかもしれない. （ A.Gyoja [20] 参照. ）なお D.Mumford, J.Fogarty and F.Kirwan [1], Appendix に, この項に関する解説がある. この Appendix は（概均質ベクトル空間の理論と限らず）不変式論全般についての非常にすぐれた解説である.

裏返し変換で b-関数, 関数等式, ゼータ関数などがどう変化するかという問題は重要である. 一変数の場合, b-関数の裏返し変換に関する挙動は Shintani の公式として知られ T.Kimura [6] にある（証明は T.Kimura [1] を見よ). Shintani 氏は, また裏返し変換の下での関数等式の挙動も計算したが, 証明に若干のギャップがあった. F.Sato and H.Ochiai [1] は, そのギャップを埋め, 標数 0 の任意の局所体上の多変数ゼータ関数に拡張した形での結果を与えている. また, そこでは, b-関数の Shintani の公式も多変数版で得られている. p-進関数等式については Y.Kajima [1] もある. p-進局所ゼータ関数の裏返し変換での挙動は J.Igusa [19], 大域的ゼータ関数の場合は F.Sato [15] で調べられている. 有限体上の場合は A.Gyoja によって調べられ, J.Denef and A.Gyoja [1] で証明された関数等式の具体形を予想するために役立った (A.Gyoja [9]).

2.7 ゼータ関数の応用

(i) 数論的量の漸近評価数論的量の評価を行うとき, その量の母関数として得られる Dirichlet 級数の解析的性質（極の位置, 留数等）が重要であることはよく知られている (Tauber 型定理） . 特に関数等式を満足する場合には鋭い評価が得られる事が多い（ E.Landau [1], K.Chandrasekhran and Raghavan Narasimhan [1], M.Sato and T.Shintani [1] ） . この種の応用についてはまず T.Shintani [1], [2] を見よ. B.Datskovski and D.J.Wright [2] は, Tauber 型の議論を adelic ゼータ関数に適合する形に発展させ, GL(2) の 3 階対称テンソル表現に付随する adelic なゼータ関数の性質から代数体の 3 次拡大体の分布についての情報を引きだした. ついで D.J.Wright and A.Yukie [1] は, このケースを含むいくつかの場合に, 概均質ベクトル空間の開軌道の有理点を作用する群の有理点のなす群で軌道分解したとき, 各有理軌道が定義体の拡大体を parametrize することを見出している（関連する文献として Ikai [1] もある） . したがって, このような空間のゼータ関数の解析的性質から代数体の分布についてのさらなる情報が得られる望みがあり, B.Datskovski, D.J.Wright, A.Yukie による研究が進んでいる.

Order の中の ideal の分布を扱った C.J.Bushnell and I.Reiner [1] も概均質ベクトル空間の立場から見た方が見通しがよいと思われる.

(ii) ゼータ関数の特殊値と保型形式の次元公式 A.Selberg [1] において, いわゆる Selberg 跡公式によって保型形式の次元を計算すると離散部分群の parabolic 共役類の寄与が Dirichlet 級数の特殊値で表されるという観察が述べられている. ここに現れる Dirichlet 級数は大雑把にいうと H.Rubenthaler [6], [7] 等で調べられている parabolic type の概均質ベクトル空間に付随する Hurwitz 型のゼータ関数である. このことは rank 1 の場合に W.Hoffmann [1] で一般的に示された（ M.S.Osborne and G.Warner [1], [2], G.Warner [1], W.Hoffmann [2] も参照） . rank ≥ 2 の場合には Y.Morita [1], T.Shintani [2], T.Arakawa [2], [4], K.Hashimoto [1], [2] がある. A.Murase and T.Sugano

[1] では, T.Shintani [2] の結果を Jacobi 形式の場合に拡張し, その際, 前述した prehomogeneous affine space のゼータ関数を導入した.

次元公式とゼータ関数の特殊値との関係は, N.Kawanaka [4] における有限 reductive 群の指標公式と概均質ベクトル空間の Gauss 和との関係に極めて類似していることを注意しておく（「表現論との関係」の (2) 項参照） .

このような応用に鑑みると, ゼータ関数の特殊値の具体的な計算法を確立することは重要である. 総実代数体のゼータ関数の特殊値を contour integral を用いて計算する T.Shintani [3] による方法は, A.Kurihara[1], I.Satake [1] によって cone のゼータ関数へと拡張された. T.Arakawa [5] も Shintani の方法の拡張を扱っている. 一方, ゼータ関数を既知のより基本的なゼータ関数を用いて表わすことができれば, 特殊値の計算が易しくなることは明らかである. このようなアプローチは, 従来, 全行列環のゼータ関数などを除いて困難だと考えられていたが, T.Ibukiyama and H.Saito [1] は T.Shintani [2] で扱われた対称行列の空間のゼータ関数について explicit な計算を実行し, Riemann ゼータ関数と半整数 weight の Eisenstein 級数の Mellin 変換を用いた明示公式を得た. その結果として Siegel modular form の空間の次元公式の完全に explicit な形が予想できる. Ibukiyama-Saito の仕事を通じて, 概均質ベクトル空間のゼータ関数は, かなり多くの場合に Riemann ゼータ関数, Dirichlet の L 関数, より一般には Eisenstein 級数を用いて表わされるであろうという見通しが生まれた.

保型形式の次元公式を Selberg 跡公式を用いて表わすとき, cusp からの寄与としてゼータ関数の特殊値が現われるが, 同じ次元公式を Riemann-Roch の定理を用いて計算すると cusp に関連する幾何的不変量が現われる. こうしてゼータ関数の特殊値と各種の幾何的不変量との関係が問題となる（いわゆる Hirzebruch 予想とその一般化） . これ等の話題については M.F.Atiyah, H.Donnelly and I.M.Singer [1], [2], W.Müller [1], [2], I.Satake [2], [3], I.Satake and S.Ogata [1], S.Ogata [1] を見よ.

2.8 表現論との関係

(1) 放物型部分群の巾零根基が可換であれば, その上に Levi 部分群が概均質に作用する. このような放物型部分群からの誘導表現は比較的容易に詳しく調べることができるが, ここに概均質ベクトル空間の理論（特に b-関数）が implicit/explicit に関わる. intertwining operator については B.D.Boe [1]. 既約性については S.Suga [1]. ユニタリー性については T.Enright, R.Howe and N.Wallach [1], H.Yamada [1] などを見よ. より一般に semi-invariant の b-関数が表現の種々の性質をコントロールしていることがわかりつつある. M.Kashiwara [5], A.Gyoja [12], [13] を見よ.

(2) 有限 reductive 群の指標表を決定するという問題と, 概均質ベクトル空間の理論は密接な関係がある. generalized Gelfand-Graev 表現の理論については, N.Kawanaka [1], [3], [4] を見よ. character sheaf については G.Lusztig [4] を見よ. T.Shoji [1] は後者についての概説である. T.Shoji [2] は指標表の決定の現状の総合報告.

(3) Dynkin-Kostant 理論により, 半単純リー環の巾零軌道の研究が概均質ベクトル空間の研究に帰着される. 一般論については T.A.Springer and R.Steinberg [1] を見よ. 具体的な情報は同所, および A.G.Elašvili [1], A.V.Alekseevsky [1] を見よ. これらを用いると Dynkin-Kostant 理論で現れる概均質ベクトル空間の generic isotropy group の具体的な形がわかる. 議論の詳細については N.Kawanaka [5] を見よ. Dynkin-Kostant 理論で現われる概均質ベクトル空間の特徴付けは V.G.Kac [1], A.Gyoja [6] にある.

(4) 概均質ベクトル空間の理論, Harish-Chandra の微分方程式系, および Aomoto-Gelfand の超幾何微分方程式系は, 互いに密接に関係していることが, 幾人もの研究者により explicit/implicit に注目されている. I.M.Gelfand [2], R.Hotta [1], R.Hotta-M.Kashiwara [1], G.Lusztig [3] を見よ.

(5) 概均質性への注目は不変式論にとっても有意義である. たとえば, R.Howe [1], R.Howe and T.Umeda [1], T.Umeda [1], [2] を見よ.

(6) 群の作用する空間において, 相対不変超関数を決定するという問題は基本的である. たとえば, P.D.Methee [1], H.Rubenthaler [1], Fulvio Ricci and H.M.Stein [1] などで調べられている. M.Muro [7], [8], [11] は, 主として既約正則被約概均質ベクトル空間を対象に microlocal calculus を方法として, この問題を調べている.